OpenStudy (anonymous):

26 gallons of gas cost $80.08. what is the unit rate for 1 gallon of gas?

OpenStudy (triciaal):

26 *1gallon = $80,08 for 1 gallon divide by 26 keep the balance of the equation divide the other side by 26 also 26*1/26 = 80.08 /26= 1 gallon = $?